OpenStudy (anonymous):

Please help. |-4b-8|+|-1-b^{2}|+2b^{3} ; b=-2

OpenStudy (anonymous):

substitute b directly. :) |-4(-2)-8|+|-1-(-2)^{2}|+2(-2)^{3}=? |8-8|+|-1-4|+2(-8)=? 0 -5 -16=? -21 :)

OpenStudy (anonymous):

Thank you!!